OpenStudy (anonymous):

Henry Devine bought a new dishwasher for $320. He paid $20 down and made 10 monthly payments of $34. What actual yearly interest rate did Henry pay? A. 29.09 B. 68.75 C. 34.38 D. 14.55

OpenStudy (anonymous):

its Down Payment= $20 10 Payments=10x34, =340 Total paid=340+20=$360 Cost price was........$320 Interest=.................._____40 The formula for Simple Interest is i=Prt, where i=annual interest rate P=starting amount r= interest rate, as a percent t=time, in years From i=Prt, r=i/Pt i=40, P=320, t=10/12 of a year r=40/320(10/12) r=40/266.67 r=0.1499981 r=15%

OpenStudy (anonymous):

@KaelenDrake did you get what i just did.?

OpenStudy (anonymous):

the answer is D.